How many liters of a 40 percent solution and a 16 percent solution must be mixed to obtain 20 liters of a 22 percent solution?
Answer by edjones(8007)

.40* x = .4x
.16* (20-x) = 3.2-.16x
.22* 20 = 4.4
.
.4x+3.2-.16x=4.4
.24x=1.2
x=5 liters of 40%
15 liters of 16%
.
Ed